If f(x)= 2(x-1)^2+3x find f(a+1)
Answer by stanbon(75887)

If f(x)= 2(x-1)^2+3x find f(a+1)
----
f(a+1) = 2(a+1-1)^2 + 3(a+1)
----
f(a+1) = 2a^2 + 3a + 3
========================
Cheers,
Stan H.